### v5

@@ Abstract.
Generative AI is moving rapidly from research into real world deployment across sectors, which elevates the need for responsible development, deployment, evaluation, and governance. To address this pressing challenge, in this study, we synthesize the landscape of responsible generative AI across methods, benchmarks, and policies, and connects governance expectations to concrete engineering practice. We follow a prespecified search and screening protocol focused on post-ChatGPT era with selective inclusion of foundational work for definitions, and we conduct a narrative and thematic synthesis. Three findings emerge; First, benchmark and practice coverage is dense for bias and toxicity but relatively sparse for privacy and provenance, deepfake and media integrity risk, and system level failure in tool using and agentic settings. Second, many evaluations remain static and task local, which limits evidence portability for audit and lifecycle assurance. Third, documentation and metric validity are inconsistent, which complicates comparison across releases and domains. We outline a research and practice agenda that prioritizes adaptive and multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration and uncertainty reporting, versioned and documented artifacts, and continuous monitoring. Limitations include reliance on public artifacts and the focus period, which may under represent capabilities reported later. The survey offers a path to align development and evaluation with governance needs and to support safe, transparent, and accountable deployment across domains.
@@ 1. Introduction
The rapid advancement of Artificial Intelligence (AI) has transformed it from a theoretical construct into a pervasive technology that shapes modern society. AI systems are embedded in critical decision making across healthcare (Kaur et al., 2022), finance (Cao, 2022a), transportation (Abduljabbar et al., 2019), manufacturing industry (Jiang et al., 2023), education (Zhai et al., 2021), and others (Hadi et al., 2023). As these systems influence outcomes that affect individuals, communities, and economies, the need for ethical, transparent, and accountable deployment has become urgent (Liu et al., 2025b).
Responsible AI (RAI) (Stryker, 2024) has emerged as a central area of research and practice to address these challenges. RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020). By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance (Raza et al., 2024a), RAI seeks to realize the societal benefits of AI while minimizing potential harms (Lu and Wu, 2025).
@@
There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025b). Nevertheless, notable gaps remain in alignment with human values and in operational practice. The academic literature often separates conceptual frameworks (Lu et al., 2024b; Saeed and Omlin, 2023; Minh et al., 2022; Mehrabi et al., 2021a) from practical implementations (Huang et al., 2024a; Dwivedi et al., 2023; Caton and Haas, 2024), as summarized in Table 1. Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts. The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (nis, 2024), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (Systems, 2021). This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
The emergence of powerful generative AI models, including conversational and multimodal systems, also require a reassessment of responsible practices (Mhlanga, 2023). Generative AI (GenAI) is transforming products and workflows across domains (Kenthapadi et al., 2023). Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization wide RAI strategy (Capgemini Research Institute, 2024). This imbalance underscores the need to ensure that development and deployment align with ethical standards, fairness, privacy, security, and societal values (Pan and Xu, 2025).
Problem statement. Current RAI guidance for GenAI is fragmented across principles, processes, and tools (as shown in Table 1) and there is less focus on governance-aligned metrics and AI-ready testbeds that enable comparable, auditable assessment across the lifecycle; especially for GenAI systems. As a result, organizations cannot reliably connect regulatory obligations to concrete engineering controls or produce portable evidence for audits and post-market monitoring.
This survey identifies a central problem. Current RAI guidance is fragmented across principles, processes, and tools, and it lacks consistent metrics and AI-ready testbeds that enable comparable assessment across the lifecycle of GenAI systems. Existing surveys (Table 1) and frameworks rarely integrate governance requirements with actionable engineering practices and with systematic evaluation protocols for generative models, including multimodal systems and agentic pipelines.
@@
    We present a unified synthesis that maps governance principles to concrete engineering controls and assurance activities across the AI lifecycle.
@@
    We present a governance-aligned comparative rubric (C1–C10) with normalized scores and a policy crosswalk (NIST AI RMF, EU AI Act, ISO/IEC 42001) linking benchmarks to audit evidence.
@@
    We also present a practice roadmap with lifecycle KPIs, explainability guidance for GenAI works, and AI-ready testbeds to support continuous, auditable evaluation.
[…]
Responsible GenAI
Challenges
Technical
Data-related: representativeness(Raza et al., 2024c), quality(Raza et al., 2025b), ethics(Gebru et al., 2021)
Model & system-related: bias(Pessach and Shmueli, 2022), privacy(Dwork, 2006), misinformation(Wang et al., 2023a)
Regulatory & Sociotechnical
IP(Kenthapadi et al., 2023), alignment(Bai et al., 2022), explainability(Dwivedi et al., 2023), sustainability(Khan et al., 2025), dual use(Weidinger et al., 2022)
Strategies
Safety benchmarks
HELM(Liang et al., 2023), SafetyBench(Zhang et al., 2024), Rainbow Teaming(Samvelyan et al., 2024), ALERT(Tedeschi et al., 2024)
Key Performance Indicators
Statistical parity difference(Kun, 2015), disparate impact(Barocas and Selbst, 2016), emissions(Khan et al., 2025)
Explainability methods
SHAP(Lundberg and Lee, 2017), LIME(Ribeiro et al., 2016a), integrated gradients(Arrieta et al., 2020), attention maps(Jain and Wallace, 2019; Serrano and Smith, 2019)
AI-ready testbeds
AI4EU(Foundation, 2024), IEEE Testbed(Foundation, 2024), logging(nis, 2024), red-teaming(Samvelyan et al., 2024; Tedeschi et al., 2024)
Governance frameworks
NIST AI RMF(nis, 2024), ISO/IEC 42001(Systems, 2021), EU AI Act(Outeda, 2024), OECD principles(Dignum, 2019)
Technical toolkits
Model cards(Liang et al., 2023), datasets datasheets(Gebru et al., 2021) , C2PA provenance standards(Das et al., 2025)
Applications
Healthcare
Concerns: Privacy(Dwork, 2006), hallucination(Wang et al., 2023c), bias in diagnostics(Kaur et al., 2022)
Solutions: RAG(Lewis et al., 2020), HITL(Weidinger et al., 2021), Differential Privacy(Dwork, 2006), FDA Alignment(Laux et al., 2024)
Finance
Concerns: Discriminatory lending(Barocas and Selbst, 2016), security(Staab et al., 2024), manipulation(Wang et al., 2023a)
Solutions: Fairness audits(Pessach and Shmueli, 2022), SHAP/LIME(Lundberg and Lee, 2017; Ribeiro et al., 2016a), Basel III/SEC(Slattery et al., 2024)
Education
Concerns: Cultural insensitivity(Al-kfairy et al., 2024), privacy(Dwork, 2006), skills evasion(Lin et al., 2024)
Solutions: Fairness audits(Pessach and Shmueli, 2022), Homomorphic encryption(Bashir et al., 2024), ethics(Dignum, 2019)
Arts & Entertainment
Concerns: Copyright(Kenthapadi et al., 2023; Freeman et al., 2024), job displacement(Pan and Xu, 2025), deepfakes(Shi et al., 2024)
Solutions: Artist collaboration(Freeman et al., 2024), Deepfake detection(Shi et al., 2024)
Agriculture
Concerns: hallucination(Raza et al., 2025d), data ownership(Mittal et al., 2024), digital divide(Al-kfairy et al., 2024)
Solutions: Geo-aware RAG(Lewis et al., 2020), Data contracts(Mittal et al., 2024), HITL(Weidinger et al., 2021)
Defense
Concerns: Autonomous weapons(Weidinger et al., 2022), opacity(Dwivedi et al., 2023), adversarial attacks(Zhang et al., 2024)
Solutions: Human Control(Weidinger et al., 2021), Red teaming(Samvelyan et al., 2024), NATO Principles(Outeda, 2024)
Figure 1. Responsible AI landscape across challenges, strategies, and applications.
This review is a structured literature review of theory, applications, and policy for Responsible AI with emphasis on GenAI. It follows a prespecified search followed by a narrative and thematic synthesis. Themes covered include explainability (XAI), fairness/bias, privacy/security, robustness/safety, governance/compliance, benchmarking and KPIs, and sectoral applications. We then operationalize the agenda via a governance-aligned evaluation rubric with a policy crosswalk and a stakeholder-focused practice roadmap with lifecycle KPIs and AI-ready testbeds. A mindmap of the paper is shown in Figure 1.
Table 1. Comparison of literature by venue, year, topic, and coverage across computer science theory, applications, policy, and generative AI. Checkmarks (✓) indicate coverage.
Paper Venue Year Topic Coverage
CS Theory Application Policy GenAI
(Schiff et al., 2020) arXiv 2020 Responsible AI
(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness
(Hu et al., 2021) ACM CSUR 2021 AI Safety
(Trocin et al., 2023) Springer ISF 2021 Responsible AI
(Minh et al., 2022) Springer AI Review 2022 Explainable AI
(Huang et al., 2024a) Springer AI Review 2023 Safety
(Saeed and Omlin, 2023) Elsevier KBS 2023 Explainable AI
(Dwivedi et al., 2023) ACM CSUR 2023 Explainable AI
(Nauta et al., 2023) ACM CSUR 2023 Bias and Fairness
(Huang et al., 2023a) IEEE TAI 2023 AI Ethics
(Díaz-Rodríguez et al., 2023) Elsevier Inf. Fusion 2023 Responsible AI
(Mhlanga, 2023) Springer Nature 2023 Responsible AI
(Lu et al., 2024b) ACM CSUR 2024 Responsible AI
(Leschanowsky et al., 2024) Elsevier CHB 2024 Privacy and Safety
(Caton and Haas, 2024) ACM CSUR 2024 Bias and Fairness
(Li and Goel, 2024) Springer ISF 2024 Accountability and Transparency
(Slattery et al., 2024) arXiv 2024 Responsible AI
(Jedličková, 2024) Springer AI and Society 2024 Responsible AI
(Al-kfairy et al., 2024) MDPI Informatics 2024 AI Ethics
(Sadek et al., 2025) AI and Society 2024 Challenges of Responsible AI in Practice
(Woodgate and Ajmeri, 2024) ACM CSUR 2024 Responsible AI
(Reuel et al., 2025) ACM FAccT 2025 Responsible AI (Global maturity model)
(Gadekallu et al., 2025) arXiv 2025 Best Practices of Responsible AI
(Byrne, 2025) ASEE Peer 2025 U.S. Responsible AI Policy
(Sadek et al., 2025) Springer AI & Society 2025 Responsible AI
(Baldassarre et al., 2025) IEEE TAI 2025 Responsible AI
(Ivanov, 2025) The Service Industries Journal 2025 Responsible use of AI in social science
Ours ACM CSUR 2025 Responsible AI
The rest of this paper is organized as follows. Section 2 presents the literature review methodology. Section 3 provides background on Responsible Generative AI. Section 4 reports our results, including thematic findings and benchmark assessments. Section 4 also discusses key performance indicators, explainable AI, and AI-ready testbeds, and Section 5 discuss RAI applications, challanges and RAI strategies. Section 6 presents discussion and forward thoughts, and finally, Section 7 concludes with open challenges and future research directions.
2. Literature Review Methodology
We present our literature review methodology in this section.
Objectives and Review Questions.
This article conducts a literature review of RAI in the post-GPT period and the synthesis of the findings. We address the following key questions:
  • RQ1: What are the dominant themes, risks, and controls in RAI for GenAI and multimodal systems since late 2022?
@@ 2. Literature Review Methodology
Scope and Protocol.
The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017; Keele et al., 2007), and we followed a structured literature review rather than a traditional systematic mapping study or an empirical meta analysis. We mainly focused on literature published from November 2022 through August 2025 to capture rapid methodological and governance developments associated with contemporary generative models. Foundational work published before 2022 was selectively included to define terms, situate concepts, and trace relevant intellectual lineages. We conducted a narrative and thematic synthesis aligned with the research questions and did not perform a meta analysis because of heterogeneity in study designs, measures, and outcomes.
Sources and Search Strategy.
Sources included the ACM Digital Library, IEEE Xplore, SpringerLink, ScienceDirect, arXiv, DBLP, Google Scholar, and Web of Science. To mitigate coverage bias, we complemented database queries with backward and forward snowballing from seed works identified as influential during pilot screening. Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity, open science), with syntax adapted to each database (full strings in Appendix A).
Eligibility and Selection.
We included English-language works within the target time window that addressed RAI in the context of GenAI or multimodal systems and appeared in peer-reviewed venues (journals or conferences), together with a small number of high-impact governance reports or widely cited preprints when they materially shaped discourse. We excluded low-credibility grey literature, purely opinionated pieces without substantive argumentation, duplicates, and items whose primary contribution lay outside RAI (e.g., domain applications without responsibility, safety, or governance components). Although our emphasis was the post-GPT era, we allowed earlier seminal contributions when necessary for definitional clarity or historical continuity.
Screening Process and Data Analysis.
Titles/abstracts were screened against the inclusion criteria, followed by full-text review for items deemed potentially eligible; screening was performed by the first author with a second pass on borderline cases. For each included record we recorded the citation key, year, publication type (journal/conference/preprint), title, and (where available) author-provided keywords. We then applied a lightweight text-mining pipeline to infer themes (e.g., fairness/bias, explainability, governance/compliance, privacy/security, robustness, safety, agentic, multimodal) from titles and keywords (abstracts when available), and aggregated counts by year and publication type. Tables 2 provide basic traceability from literature to claims.
Theme Citation Year Focus
Safety/Benchmarks (Liu et al., 2024) 2024 Introduces a multimodal benchmark to assess safety risks in vision–language models.
Safety/Benchmarks (Zhang et al., 2024) 2024 Proposes a general safety benchmark for large language models across risk categories.
Safety/Benchmarks (Gehman et al., 2020) 2020 Measures toxic degeneration using curated prompts to stress test language models.
Safety/Benchmarks (Li et al., 2024a) 2024 Presents a hierarchical and comprehensive safety benchmark for language models.
Safety/Benchmarks (Mazeika et al., 2024) 2024 Standardizes evaluation for red teaming and refusal robustness in language models.
Safety/Benchmarks (Samvelyan et al., 2024) 2024 Generates diverse adversarial prompts for automated safety stress testing.
Safety/Benchmarks (Wang et al., 2023a) 2023 Provides a broad audit of trustworthiness, including safety and robustness, for GPT models.
Safety/Benchmarks (Tedeschi et al., 2024) 2024 Benchmarks red teaming effectiveness for assessing model safety responses.
Fairness/Governance (Mittal et al., 2024) 2024 Outlines dataset design practices aligned with fairness, privacy, and regulatory norms.
Fairness/Governance (Rodolfa et al., 2021) 2021 Finds minimal fairness and accuracy tradeoff in public policy machine learning tasks.
Fairness/Governance (Chouldechova, 2017) 2017 Analyzes disparate impact and fairness criteria in recidivism risk prediction.
Fairness/Governance (Barocas and Selbst, 2016) 2016 Examines how big data practices can produce disparate impact and legal risk.
Fairness/Governance (Pessach and Shmueli, 2022) 2022 Surveys algorithmic fairness methods and evaluation across application domains.
Fairness/Governance (Kaur et al., 2022) 2022 Reviews trustworthy AI principles and technical enablers for responsible deployment.
Fairness/Governance (Dignum, 2019) 2019 Presents a framework for responsible AI grounded in ethics and governance.
Fairness/Governance (Laux et al., 2024) 2024 Discusses trustworthiness concepts in relation to compliance under the EU AI Act.
Explainability/XAI (Ribeiro et al., 2016b) 2016 Introduces LIME for model agnostic local explanations of predictions.
Explainability/XAI (Selvaraju et al., 2017) 2017 Proposes Grad-CAM for visual explanations in convolutional neural networks.
Explainability/XAI (Arrieta et al., 2020) 2020 Synthesizes concepts, taxonomies, and challenges in explainable artificial intelligence.
Explainability/XAI (Dwivedi et al., 2023) 2023 Reviews core ideas and techniques for explainable AI with practical guidance.
Explainability/XAI (Saeed and Omlin, 2023) 2023 Provides a meta survey of XAI challenges and future research directions.
Explainability/XAI (van der Waa et al., 2021) 2021 Compares rule based and example based explanations on fidelity and user value.
Explainability/XAI (Mosca et al., 2022) 2022 Reviews SHAP based explanation methods for natural language processing.
Explainability/XAI (Nauta et al., 2023) 2023 Systematizes evaluation methods for explanations with quantitative protocols.
Privacy/Security (Dwork, 2006) 2006 Establishes differential privacy as a formal framework for privacy protection.
Privacy/Security (McMahan et al., 2017) 2017 Introduces federated learning for decentralized training with privacy benefits.
Privacy/Security (Nasr et al., 2023) 2023 Demonstrates large scale extraction of memorized training data from language models.
Privacy/Security (Mattern et al., 2023) 2023 Proposes membership inference against language models via neighborhood comparison.
Privacy/Security (He et al., 2025) 2025 Presents label only membership inference attacks tailored to pre trained language models.
Privacy/Security (Li et al., 2024b) 2024 Introduces a benchmark to evaluate privacy risks in language models at multiple levels.
Privacy/Security (Mireshghallah et al., 2024) 2024 Tests privacy leakage using contextual integrity based evaluations of language models.
Privacy/Security (Staab et al., 2024) 2024 Assesses the potential of large models for anonymization and de identification tasks.
LLM Safety (Bai et al., 2022) 2022 Aligns models using constitutional principles and self critique to reduce harmful outputs.
LLM Safety (Weidinger et al., 2021) 2021 Maps ethical and social risks from language models and proposes mitigation avenues.
LLM Safety (Huang et al., 2024b) 2024 Positions trustworthiness dimensions and evaluation needs specific to large models.
LLM Safety (Lin et al., 2024) 2024 Surveys techniques for debiasing and reducing hallucination in large models.
LLM Safety (Wang et al., 2023c) 2023 Overviews ChatGPT applications, limitations, and emerging risks and safeguards.

### v6

@@ Abstract.
Generative AI is rapidly moving from research to deployment, elevating the need for responsible development, evaluation, and governance. We conduct a PRISMA-guided review of 232 studies (November 2022–December 2025), spanning large language models, vision-language models, diffusion models, and agentic pipelines. We make four contributions: (1) the first survey bridging governance principles, technical evaluation, and domain deployment across all four system types; (2) a ten-criterion rubric (C1–C10) scoring major AI safety benchmarks on risk-surface coverage, paired with a policy crosswalk mapping benchmarks to regulatory requirements; (3) twelve lifecycle KPIs, explainability guidance for foundation models, and a testbed catalogue; and (4) domain-specific analysis across healthcare, finance, education, arts, agriculture, and defense. Three findings emerge: benchmark coverage is dense for bias and toxicity but sparse for privacy, provenance, deepfakes, and system-level failures in agentic settings; evaluations remain largely static and task-local, limiting audit portability; and inconsistent documentation complicates cross-release comparison. We outline a research agenda prioritizing adaptive multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration reporting, versioned artifacts, and continuous monitoring. This survey offers a structured path to align generative AI evaluation with governance needs for safe and accountable deployment.
@@ 1. Introduction
Generative AI (GenAI) systems, including large language models (LLMs), vision language models (VLMs), diffusion models, and emerging agentic pipelines, are rapidly transitioning from research prototypes to production deployments that generate text, images, code, and autonomous actions across healthcare, finance, education, and other sectors (Hadi et al., 2023; Hajkowicz et al., 2023). This accelerated deployment heightens the need for responsible development practices that address GenAI-specific risks, including hallucination and unverifiability, prompt injection and jailbreak vulnerabilities, data memorization and leakage, deepfake and media integrity threats, and system-level failures in tool-augmented settings.
Responsible AI (RAI) (Stryker, 2024) has emerged as a central area of research and practice to address these challenges. RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020). By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance, RAI seeks to realize the societal benefits of AI while minimizing potential harms. While these principles were initially developed for predictive systems, generative models introduce qualitatively new challenges: outputs are open-ended and difficult to enumerate, stochastic decoding produces non-deterministic behaviour, multimodal synthesis creates cross-modal failure modes, and agentic tool use extends the attack surface beyond the model itself. Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization-wide responsible AI strategy (Capgemini Research Institute, 2024). This imbalance underscores the need to align the development and deployment of these systems with ethical standards, fairness, privacy, security, and societal values.
@@
There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025). Nevertheless, notable gaps remain in alignment with human values and in operational practice. The academic literature often separates conceptual frameworks from practical implementations, as summarized in Table 1. Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts. The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (of Standards and, NIST), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (ISO/IEC, 2021b). This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
In this study, we conduct a comprehensive literature review of responsible GenAI (ResGenAI) frameworks, governance guidelines, evaluation benchmarks, and operational risk assessment methodologies. Our scope centres on the application of RAI principles to generative systems, including LLMs, VLMs, diffusion models, and agentic/RAG pipelines. While we draw on the broader RAI literature for foundational constructs (e.g., fairness, privacy, safety, transparency, accountability), our analysis targets ResGenAI specific risk mechanisms such as hallucination and unverifiability, prompt injections, jailbreaking and tool-use failures, data memorization and leakage, provenance and deepfakes or media integrity risks, intellectual property concerns, and system-level failures in GenAI applications. Despite substantial prior work, ResGenAI guidance remains fragmented across principles, processes, and tools (Table 1). Consequently, there is a pressing need to translate regulatory obligations into concrete engineering controls and to produce portable audit evidence for post-market monitoring.
{forest}
Figure 1. Responsible GenAI landscape organised into three pillars: Challenges (Section 3), Strategies (Section 4), and Applications (Section 5). Technical challenges span data, model/system, and emergent-threat layers; strategies encompass benchmarks, KPIs, explainability, testbeds, governance, and toolkits; applications map domain-specific risks and controls across six sectors.
@@
    To the best of our knowledge, this is the first review article that simultaneously spans governance principles, technical evaluation, and domain deployment for GenAI. We synthesize findings across the NIST AI RMF, EU AI Act, ISO/IEC 42001, and other instruments listed in Table 6, covering LLMs, VLMs, diffusion models, and RAG topics. As shown in Table 1, no prior survey bridges all four thematic areas (CS theory, application, policy, GenAI) and all four novelty dimensions (taxonomy, scope, evaluation, methods) simultaneously.
@@
    We introduce a ten-criterion rubric (C1–C10) that scores widely used AI safety benchmarks on risk-surface coverage and assurance quality, and a policy crosswalk (Section 4.2.1) that maps each benchmark to the governance requirements it can plausibly evidence. The closest efforts (Liu et al., 2025b; Ferdaus et al., 2026) partially address evaluation taxonomy or trust assessment, but neither links benchmarks to regulatory audit-evidence needs.
@@
    We operationalise the rubric findings into measurable KPIs (Section 4.2.4), explainability guidance tailored to foundation models (Section 4.3), and a testbed catalogue for continuous evaluation. Section 4.2 links these KPIs to specific benchmarks and audit toolkits, providing practitioners with a decision-support matrix for assembling governance-ready evidence.
  4. (4)
    We apply the RAI lens to real-world applications and domains (Section 5, demonstrating how technical advancements can be effectively aligned with governance frameworks and regulatory requirements.
Table 1. Comparison of related surveys by venue, year, topic, thematic coverage, and novelty dimensions. Checkmarks (✓) indicate coverage; (x) indicates no explicit coverage; (★) indicates partial or limited treatment. Novelty dimensions: Taxonomy = structured risk taxonomy or classification; Scope = multimodal + agentic coverage; Evaluation = governance-aligned benchmarking/rubric; Methods = actionable engineering controls or lifecycle KPIs
Paper Venue Year Topic Thematic Coverage Novelty Dimensions
CS Theory
Application
Policy
GenAI
Taxonomy
Scope
Evaluation
Methods
(Schiff et al., 2020) arXiv 2020 Responsible AI x x x x x
(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness x x x x x x
(Hu et al., 2021) ACM CSUR 2021 AI Safety x x x x x x
(Trocin et al., 2023) Springer ISF 2021 Responsible AI x x x x x
(Minh et al., 2022) Springer AI Rev. 2022 Explainable AI x x x x x
(Huang et al., 2024a) Springer AI Rev. 2023 Safety x x x x
(Saeed and Omlin, 2023) Elsevier KBS 2023 Explainable AI x x x x x x
(Dwivedi et al., 2023) ACM CSUR 2023 Explainable AI x x x x x x
(Nauta et al., 2023) ACM CSUR 2023 Bias and Fairness x x x x x
(Huang et al., 2023a) IEEE TAI 2023 AI Ethics x x x x x
(Díaz-Rodríguez et al., 2023) Elsevier Inf. Fusion 2023 Responsible AI x x x
(Mhlanga, 2023) Springer Nature 2023 Responsible AI x x x x
(Lu et al., 2024b) ACM CSUR 2024 Responsible AI x x x x x
(Leschanowsky et al., 2024) Elsevier CHB 2024 Privacy and Safety x x x x x
(Caton and Haas, 2024) ACM CSUR 2024 Bias and Fairness x x x x x
(Li and Goel, 2024) Springer ISF 2024 Accountability x x x x x
(Slattery et al., 2024) arXiv 2024 Responsible AI x x x x x x
(Jedličková, 2024) Springer AI & Soc. 2024 Responsible AI x x x x x x
(Al-kfairy et al., 2024) MDPI Informatics 2024 AI Ethics x x x x x
(Sadek et al., 2025) AI and Society 2024 RAI in Practice x x x x x
(Woodgate and Ajmeri, 2024) ACM CSUR 2024 Responsible AI x x x x x
(Reuel et al., 2025) ACM FAccT 2025 RAI (Global maturity) x x
(Gadekallu et al., 2025) arXiv 2025 Best Practices of RAI x x x x x
(Byrne, 2025) ASEE Peer 2025 U.S. RAI Policy x x x x x x x
(Sadek et al., 2025) Springer AI & Soc. 2025 Responsible AI x x x x x
(Baldassarre et al., 2025) IEEE TAI 2025 Responsible AI x x x x x
(Ivanov, 2025) Serv. Ind. J. 2025 RAI in Social Sci. x x x x x x
(Yu et al., 2025) ACM KDD 2025 Trustworthy LLM Agents x x x
(Dong et al., 2025) Springer AI Rev. 2025 LLM Safeguards x x
(Ferdaus et al., 2026) ACM CSUR 2025 Trustworthy LLMs x x x x
(Wang et al., 2025) arXiv 2025 Reasoning Model Safety x x x
(He et al., 2025b) ACM CSUR 2025 LLM Agent Security x x x x
(Papagiannidis et al., 2025) Elsevier JSIS 2025 RAI Governance x x x x x
(Liu et al., 2025b) arXiv 2025 LLM Safety Evaluation x x x x
Ours ACM CSUR 2026 Responsible GenAI
Column definitions CS Theory: conceptual frameworks, taxonomies, or algorithmic contributions. Application: domain-specific deployment or case studies. Policy: governance instruments, regulatory analysis, or compliance guidance. GenAI: explicit focus on generative models (LLMs, VLMs, diffusion, agentic). Taxonomy: structured risk classification beyond a flat listing of concerns. Scope: coverage of multimodal systems and agentic/tool-using pipelines. Evaluation: governance-aligned benchmarking with a scoring rubric or crosswalk. Methods: actionable engineering controls, lifecycle KPIs, or testbed guidance. ★ = partial or limited treatment.
The rest of this paper is organized as follows. Section 2 presents the literature review methodology. Section 3 provides background on ResGenAI. Section 4 reports our results, including thematic findings, benchmark assessments, key performance indicators, explainable AI, and AI-ready testbeds. Section 5 discusses ResGenAI applications, challenges, and strategies. Section 6 presents discussion and forward thoughts, and finally, Section 7 concludes with open challenges and future research directions. The taxonomy presented in Figure 1 provides a structural blueprint for this survey, establishing the thematic connections between high-level technical challenges and domain-specific applications.
2. Methodology
This review adheres to the PRISMA guidelines (Page et al., 2021) to ensure a transparent, reproducible, and rigorous synthesis of evidence. This study follows established methodological recommendations for performing systematic reviews within the computer science community (Al-Selwi et al., 2024; Boland et al., 2017a; Keele et al., 2007). To ensure comprehensive and consistent reporting, the structural framework was modelled after previous extensive surveys in the domain. The mapping procedure utilized in this review is organized into four interconnected phases that covering objectives (Section 2.1), search strategy (Section 2.2), eligibility criteria (Section 2.3), and study selection (Section 2.4) which are visualized in the methodological flow (Fig. 2) and detailed throughout Section 2. For a more exhaustive exploration of these mapping and reporting protocols, foundational works in the field (Boland et al., 2017a; Al-Selwi et al., 2024; Keele et al., 2007; Boland et al., 2017b) offer additional guidance.
Figure 2. The systematic literature mapping process utilized in this study
2.1. Research Questions Formulation
This study conducts a literature review of responsible generative AI (ResGenAI) in the post-GPT period and the synthesis of the findings. We address the following key questions:
  • RQ1: What are the dominant themes, risks, and controls in ResGenAI and multimodal systems since late 2022?
    Motivation: The rapid transition of GenAI from research to production has outpaced traditional risk taxonomies, necessitating a contemporary synthesis of how multimodal and agentic systems introduce qualitatively new failure modes beyond predictive ML.
@@ 2.1. Research Questions Formulation
    Motivation: Identifying the “Evaluation Gap” is critical for determining if current safety suites adequately model the adaptive, iterative pressure applied by real-world adversaries versus merely serving as static, keyword-based refusal scripts.
@@
    Motivation: A structural deficit in evidence portability hinders institutional trust; this question seeks to evaluate if safety claims are grounded in verifiable technical artifacts or subjective, unvalidated appraisals.
@@
    Motivation: Principles without practice lead to “safetywashing.” We aim to bridge the gap between regulatory mandates like the EU AI Act and concrete engineering controls, providing an implementation path for verifiable stewardship.
2.2. Evidence Acquisition and Search Strategy
A comprehensive search was conducted across five major academic databases and registers: Scopus ($n=1502$), Web of Science ($n=2798$), IEEE Xplore ($n=279$), ACM Digital Library ($n=656$), and arXiv ($n=78$). To capture the surge in GenAI research following the release of ChatGPT, the search window was limited to publications from November 2022 through December 2025. The search strings utilized an intersection of terms related to Generative AI (e.g., “Large Language Model”, “Foundation Model”, “LLM”) and Responsibility facets (e.g., “AI Safety”, “Trustworthy AI”, “Alignment”, “Hallucination mitigation”). Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity), with syntax adapted to each database (full strings in Appendix A.1). Syntax was adapted for each database to support field-specific constraints (Title/Abstract/Keywords), and proximity operators were employed where available to enhance precision. To mitigate coverage bias, database queries were complemented with forward and backward snowballing from seed works identified as influential during pilot screening. The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017a; Keele et al., 2007)
2.3. Eligibility Criteria
To ensure methodological rigor and relevance to the research questions (RQ1–RQ4), we applied the following criteria:
  • Inclusion Criteria: (1) Peer-reviewed journals, conference proceedings, or high-impact preprints materially shaping RAI discourse; (2) English-language works published within the target 2022–2025 window; and (3) Studies proposing technical frameworks, evaluation metrics, or empirical risk assessments for generative or multimodal systems.
  • Exclusion Criteria: (1) Purely opinionated pieces or editorials without substantive technical argumentation; (2) Short papers or reports lacking sufficient methodological depth; (3) Items whose primary contribution lay outside RAI (e.g., domain applications without safety or governance components); and (4) Redundant duplicates or inaccessible records.
Figure 3. PRISMA flowchart of the systematic literature identification, screening, and inclusion process
2.4. Study Selection and Screening
As detailed in the PRISMA flow diagram (Fig. 3), the initial search identified 5,313 records. Selection was conducted by three independent reviewers to minimize bias. Following automated deduplication ($n=2125$) and the removal of records marked as ineligible by automation tools ($n=609$) or other technical reasons ($n=53$), 2,562 unique records remained for screening.
Titles and abstracts were screened against the eligibility criteria, resulting in the exclusion of 2,153 items. Subsequently, 372 reports were sought for full-text retrieval and assessed for eligibility. Reports were excluded if they failed to meet quality and technical thresholds ($n=140$) or could not be retrieved ($n=37$). This rigorous multi-reviewer process yielded a final synthesis set of 232 studies. Data extraction followed a thematic pipeline to categorize findings into risks, controls, and evaluation facets as defined by the research objectives.